# Bayesian RG Flow in Neural Network Field Theories

### v1

@@ 3.2 An Information Theoretic form of Wilsonian Renormalization
Fig. 4: BRG is an information geometric coarse graining scheme that realizes a family of parameter distributions $\pi_{\Lambda}$ indexed by a distinguishability scale $\Lambda$.
@@
From a statistical inference perspective, the Fisher metric plays a rather central role. In this paper, we will concentrate on the fact that the Fisher metric provides an automatic quantification of the relevance of parameters appearing in given model [95]. In particular, the diagonal elements [^13] Strictly speaking it would be more geometrically meaningful to diagonalize the Fisher metric so that the following statements are basis independent. With that being said, diagonalization is very computationally costly and in a model with a sparse Fisher metric the diagonal components in a chosen basis are a good proxy for the eigenvalues [93]. of the Fisher matrix, $\lambda_{A}\equiv\mathcal{I}_{AA}(\theta)$, may be regarded as quantifying the sensitivity of the $A^{th}$ parameter to the acquisition of new data conditional on the data generating model having parameter estimate $\theta$. In this context the following terminology is often utilized to categorize parameters: We say that the $A^{th}$ parameter is more stiff than the $B^{th}$ parameter if $\lambda_{A}/\lambda_{B}>1$. Conversely, if $\lambda_{A}/\lambda_{B}<1$, we say that the $A^{th}$ parameter is more sloppy than the $B^{th}$ parameter.
In general, a stiff parameter is one which can be reliably fit with a reasonable amount of data. By contrast, sloppy parameters tend to require extremely large or precise data in order to fit. For this reason, a model which possesses a large number of sloppy parameters retains a high degree of degeneracy in the sense that these parameters may be varied largely without impacting the performance of the model dramatically. [^14] That is, large changes in the actual sloppy parameter value $\theta^{A}$ correspond to comparatively short distances as measured by the information metric, and in turn leave the model effectively unchanged in the relative entropy sense. Information geometrically, the diagonal element $\lambda_{A}$ is a proxy for the radius of the model space in the $A^{th}$ parameter direction. Thus, one may qualitatively regard the model space as consisting of some number of “macroscopic” dimensions corresponding to the stiff parameters, and some number of highly compact dimensions corresponding to the sloppy ones.
@@ 3.3 Information Shell Renormalization for general inference problems and NNFTs
Fig. 5: Applying NNFT to a family of Bayesian renormalized parameter distributions realizes a one parameter family of Euclidean field theories which we interpret as an information theoretically coarse grained flow in the space of NNFTs.
@@
Fig. 6: Schematic picture of information geometry interpretation of renormalization, where the top arrow represents the flow induced in model space by training, and the lower arrow the flow induced by Bayesian renormalization.
@@ 4.1 Bayesian RG for NNGPs
To proceed we make a simplifying assumption that the terminal posterior $\pi^{(L)}(b_{L},w_{L})$ is a Gaussian sharply peaked around the optimal gradient descent estimate $\theta_{*}\equiv(\overline{b}_{L},\overline{w}_{L})$. That is, of the form
@@ 4.1.1 BRG as mass renormalization of a generalized cos-net architecture
A GRP with kernel (76) has a density $p(\tilde{\phi})\sim\exp(-S(\tilde{\phi}))$, where
$S(\tilde{\phi})=\int d^{d}x\;\tilde{\phi}(x)(\nabla^{2}+m^{2})\tilde{\phi}(x),$ (77)
@@ 4.2.1 Implementation and network architecture
As we have mentioned, to good approximation we expect the posterior predictive distribution over network outputs to be Gaussian both at initialization and after training. Thus, in this case we should be able to recover the full distribution by computing the arithmetic mean and covariance over outputs relative to the realized parameter ensemble $\{\theta_{n}(D)\}_{n=1}^{k}$. This is a particularly simple application of the NNFT correspondence in which the first two cumulants are sufficient for performing the Edgeworth expansion [^24] In fact, practically speaking, expanding probability distributions in terms of their $n$-point functions generalises well to many realistic datasets, even for sensibly low $n$ [106].. In Section 4.1.1 we highlighted the analysis of the generalized cos-net for the striking similarity of its associated GRP to that of a free scalar field theory. However, while prior works have shown promising results from networks with cosine activations at initialization, they generally train poorly (due to their periodic nature). Therefore, we only consider NNs with ReLU activations in this numerics section. However, we highlight that the analysis of Section 4.1 is sufficiently general as to apply to any activation provided one can compute the functions (59).
@@
Fig. 7: Table of model hyperparameters.
@@ 4.2.2 Training
To mitigate confusion, we emphasize that training and renormalization of the network are disparate processes – in the current form, the network is first trained conventionally, then renormalized (in 4.2.3). While it is indeed true that the Bayesian renormalization scheme may be applied at any stage during the training, even at initialization, in practice relevant (and, inversely, irrelevant) parameters are typically not distinguished until near the end of training when the loss has stabilized [^25] This phenomena is particularly manifest in ‘grokked’ networks [107, 108]: models undergoing grokking rapidly change between memorization and generalization phases during training. In this case, the parameter relevance is not stable until post-grokking..
@@
Fig. 8: Distribution of evaluations of the test function at a characteristic validation sample fitted with a Gaussian distribution prior to renormalization, with mean approximately equal to the analytic value.
@@ 4.2.3 Renormalization
Fig. 9: Average model loss as a function of cutoff (note the logarithmic scale on the horizontal axis).
@@
Fig. 10: Distribution of ensemble realizations with changing cutoff. Figures 10(a) and 10(b) show the distribution of ensemble evaluations before the critical cutoff, whilst Figures 10(c) and 10(d) after $\Lambda_{C}$.
@@
Fig. 11: Distribution of ensemble realization characteristic statistics at fixed test data with changing cutoff.
@@ 5 Discussion
To briefly review, the main legs of the diagram coincide with BRG and NNFT. As reviewed in Section 3.1, NNFT maps pairs $(\phi_{\theta},\pi)$, comprising a NN architecture and its parameter distribution, into the space of SFTs. More explicitly, NNFT constructs an SFT with action, $S[\phi]$, whose correlation functions are equivalent to those obtained from an infinite ensemble of NN output functions. BRG (as reviewed in Section 3.2) is a map from a given parameter distribution, $\pi$, into a coarse-grained version, $\pi_{\Lambda}$. Here, coarse graining is defined as averaging over fine-grained information in parameter space as determined by the Fisher information metric. The Fisher information metric measures the level of distinguishability between models and the cutoff $\Lambda$ sets the desired scale of distinguishability. Thus, $\pi_{\Lambda}$ is a distribution in which parameters within a neighborhood of radius $\Lambda$, as measured by the Fisher information metric, are treated as indistinguishable and subsequently averaged together. Combining the main legs together, we realize the BRG-NNFT framework. We first construct a family of pairs $\{(\phi_{\theta},\pi_{\Lambda})\}_{\Lambda\in U\subseteq R^{+}}$, where each element corresponds to a NN architecture with its parameter distribution coarse grained to a distinguishability scale $\Lambda$. We then map each pair $(\phi_{\theta},\pi_{\Lambda})$ into a dual SFT with action, $S_{\Lambda}[\phi]$. The main thesis of this paper is that the family of SFT actions $\{S_{\Lambda}[\phi]\}_{\Lambda\in U\subseteq R^{+}}$ should be interpreted as an information-theoretic BRG flow.
@@
An enduring theme of this work is that of duality, both in terms of explicit analysis and in the sense of using complimentary ideas to reinforce and enrich each other. For example, on one hand, NNFT amplifies the interpretability of BRG by encoding abstract Bayesian parameter distributions in the form of more tractable SFTs. On the other hand, BRG expands the usefulness of NNFT by endowing it with an information theoretic tool for navigating between SFTs. The combined framework of BRG-NNFT straddles the boundary between physics and ML. It therefore presents exciting opportunities for advancements on either side of the physics/ML collaboration as well as at their interface [109]. In this respect, BRG-NNFT is an emerging research tool that is well situated in the landscape of physics meets ML.
@@
In closing, machine learning is in a period of explosive growth relative to its performance and applicability. For this reason it is crucial that we develop techniques for understanding and interpreting ML algorithms. NNFT [32, 33, 34, 12] and BRG [30, 31] are physics-inspired interpretability tools, and when brought together they provide the ingredients for mapping out the space of NNs and SFTs. In many ways this state of affairs closely resembles the state of quantum field theory (QFT) prior to Wilson’s seminal work on the renormalization group [35, 36]. At that time, QFT was a wildly powerful, but largely unwieldy tool. Wilsonian RG gave a methodology for organizing our thinking about QFTs which in turn has given us more control over how we can use and study them. Our hope is that the approach we have outlined in this paper can serve a similar organizing role for machine learning and the study of NNs.
@@ Acknowledgements
We wish to thank Jim Halverson for useful discussions on the NN/FT correspondence at StringData 2023. We also wish to thank David Berman for his insight and many useful discussions over the course of this work. Finally we thank Edward Hirst, Nathaniel Craig, Katy Craig, Emanuele Gendy, Ro Jefferson, and Zohar Ringel for useful discussions.

### v2

@@ 3.2 An Information Theoretic form of Wilsonian Renormalization
Figure 4: BRG is an information geometric coarse graining scheme that realizes a family of parameter distributions $\pi_{\Lambda}$ indexed by a distinguishability scale $\Lambda$.
@@
From a statistical inference perspective, the Fisher metric plays a rather central role. In this paper, we will concentrate on the fact that the Fisher metric provides an automatic quantification of the relevance of parameters appearing in given model [97]. In particular, the diagonal elements [^13] Strictly speaking it would be more geometrically meaningful to diagonalize the Fisher metric so that the following statements are basis independent. With that being said, diagonalization is very computationally costly and in a model with a sparse Fisher metric the diagonal components in a chosen basis are a good proxy for the eigenvalues [95]. of the Fisher matrix, $\lambda_{A}\equiv\mathcal{I}_{AA}(\theta)$, may be regarded as quantifying the sensitivity of the $A^{th}$ parameter to the acquisition of new data conditional on the data generating model having parameter estimate $\theta$. In this context, the following terminology is often utilized to categorize parameters: We say that the $A^{th}$ parameter is more stiff than the $B^{th}$ parameter if $\lambda_{A}/\lambda_{B}>1$. Conversely, if $\lambda_{A}/\lambda_{B}<1$, we say that the $A^{th}$ parameter is more sloppy than the $B^{th}$ parameter.
In general, a stiff parameter is one which can be reliably fit with a reasonable amount of data. By contrast, sloppy parameters tend to require extremely large or precise data in order to fit. For this reason, a model which possesses a large number of sloppy parameters retains a high degree of degeneracy in the sense that these parameters may be varied largely without impacting the performance of the model dramatically. [^14] That is, large changes in the actual sloppy parameter value $\theta^{A}$ correspond to comparatively short distances as measured by the information metric, and in turn leave the model effectively unchanged in the relative entropy sense. Information geometrically, the diagonal element $\lambda_{A}$ is a proxy for the radius of the model space in the $A^{th}$ parameter direction. Thus, one may qualitatively regard the model space as consisting of some number of “macroscopic" dimensions corresponding to the stiff parameters, and some number of highly compact dimensions corresponding to the sloppy ones.
@@ 3.3 Information Shell Renormalization for general inference problems and NNFTs
Figure 5: Applying NNFT to a family of Bayesian renormalized parameter distributions realizes a one parameter family of Euclidean field theories which we interpret as an information theoretically coarse grained flow in the space of NNFTs.
@@
Figure 6: Schematic picture of information geometry interpretation of renormalization, where the top arrow represents the flow induced in model space by training, and the lower arrow the flow induced by Bayesian renormalization. We note that the space of allowed models may converge to a different point than the true model (X). In particular, for finite data, more than one epoch of training will place artificial importance on the sampled data. This may skew the final model away from the true model. However, in the limit of infinite data and a single training epoch, the final space of allowed models and the true model will coincide.
@@ 4.1 Bayesian RG for NNGPs
To proceed we make a simplifying assumption that the terminal posterior $\pi^{(L)}(b_{L},w_{L})$ is a Gaussian sharply peaked around the optimal gradient descent estimate $\theta_{*}\equiv(\overline{b}_{L},\overline{w}_{L})$. [^20] We note that here $\overline{b}_{L},\overline{w}_{L}$ are the hyperparameters which identify the mean of the parameter distributions for biases and weights. In practical applications, $\overline{b}_{L},\overline{w}_{L}$ can be estimated via ensemble averaging. That is, of the form
@@ 4.1.1 BRG as mass renormalization of a generalized cos-net architecture
A GRP with kernel (76) has a density $p(\tilde{\phi})\sim\exp(-S[\tilde{\phi}])$, where
$S[\tilde{\phi}]=\int d^{d}x\;\tilde{\phi}(x)(\nabla^{2}+m^{2})\tilde{\phi}(x),$ (77)
@@ 4.2.1 Implementation and network architecture
As we have mentioned, to good approximation we expect the posterior predictive distribution over network outputs to be Gaussian both at initialization and after training. Thus, in this case we should be able to recover the full distribution by computing the arithmetic mean and covariance over outputs relative to the realized parameter ensemble $\{\theta_{n}(D)\}_{n=1}^{k}$. This is a particularly simple application of the NNFT correspondence in which the first two cumulants are sufficient for performing the Edgeworth expansion. [^25] In fact, practically speaking, expanding probability distributions in terms of their $n$-point functions generalises well to many realistic datasets, even for sensibly low $n$ [106]. In Section 4.1.1 we highlighted the analysis of the generalized cos-net for the striking similarity of its associated GRP to that of a free scalar field theory. However, while prior works have shown promising results from networks with cosine activations at initialization, they generally train poorly (due to their periodic nature). Therefore, we only consider NNs with ReLU activations in this numerics section. However, we highlight that the analysis of Section 4.1 is sufficiently general as to apply to any activation provided one can compute the functions (59).
@@
Figure 7: Table of model hyperparameters.
@@ 4.2.2 Training
To mitigate confusion, we emphasize that training and renormalization of the network are disparate processes – in the current form, the network is first trained conventionally, then renormalized (in 4.2.3). While it is indeed true that the Bayesian renormalization scheme may be applied at any stage during the training, even at initialization, in practice relevant (and, inversely, irrelevant) parameters are typically not distinguished until near the end of training when the loss has stabilized. [^26] This phenomena is particularly manifest in ‘grokked’ networks [107, 108]: models undergoing grokking rapidly change between memorization and generalization phases during training. In this case, the parameter relevance is not stable until post-grokking.
@@
Figure 8: Distribution of evaluations of the test function at a characteristic validation sample fitted with a Gaussian distribution prior to renormalization, with mean approximately equal to the analytic value.
@@ 4.2.3 Renormalization
Figure 9: Average model loss as a function of cutoff (note the logarithmic scale on the horizontal axis).
@@
Figure 10: Distribution of ensemble realizations with changing cutoff. Figures 10(a) and 10(b) show the distribution of ensemble evaluations before the critical cutoff, whilst Figures 10(c) and 10(d) after $\Lambda_{C}$.
@@
Figure 11: Distribution of ensemble realization characteristic statistics at fixed test data with changing cutoff.
@@ 5 Discussion
To briefly review, the main legs of the diagram coincide with BRG and NNFT. As reviewed in Section 3.1, NNFT maps pairs $(\phi_{\theta},\pi)$, comprising a NN architecture and its parameter distribution, into the space of SFTs. More explicitly, NNFT constructs an SFT with action, $S[\phi]$, whose correlation functions are equivalent to those obtained from an infinite ensemble of NN output functions. BRG (as reviewed in Section 3.2) is a map from a given parameter distribution, $\pi$, into a coarse-grained version, $\pi_{\Lambda}$. Here, coarse graining is defined as averaging over fine-grained information in parameter space as determined by the Fisher information metric. The Fisher information metric measures the level of distinguishability between models and the cutoff $\Lambda$ sets the desired scale of distinguishability. Thus, $\pi_{\Lambda}$ is a distribution in which parameters within a neighborhood of radius $\Lambda$, as measured by the Fisher information metric, are treated as indistinguishable and subsequently averaged together. Combining the main legs together, we realize the BRG-NNFT framework. We first construct a family of pairs $\{(\phi_{\theta},\pi_{\Lambda})\}_{\Lambda\in U\subseteq\mathbb{R}^{+}}$, where each element corresponds to a NN architecture with its parameter distribution coarse grained to a distinguishability scale $\Lambda$. We then map each pair $(\phi_{\theta},\pi_{\Lambda})$ into a dual SFT with action, $S_{\Lambda}[\phi]$. The main thesis of this paper is that the family of SFT actions $\{S_{\Lambda}[\phi]\}_{\Lambda\in U\subseteq\mathbb{R}^{+}}$ should be interpreted as an information-theoretic BRG flow.
@@
An enduring theme of this work is that of duality, both in terms of explicit analysis and in the sense of using complimentary ideas to reinforce and enrich each other. For example, on one hand, NNFT amplifies the interpretability of BRG by encoding abstract Bayesian parameter distributions in the form of more tractable SFTs. On the other hand, BRG expands the utility of NNFT by endowing it with an information-theoretic tool for navigating between SFTs. The combined framework of BRG-NNFT straddles the boundary between physics and ML. It therefore presents exciting opportunities for advancements on either side of the physics/ML collaboration as well as at their interface [109]. In this respect, BRG-NNFT is an emerging research tool that is well-situated in the landscape of physics meets ML.
@@
In closing, machine learning is in a period of explosive growth relative to its performance and applicability. For this reason it is crucial that we develop techniques for understanding and interpreting ML algorithms. NNFT [32, 33, 34, 12] and BRG [30, 31] are physics-inspired interpretability tools, and when brought together they provide the ingredients for mapping out the space of NNs and SFTs. In many ways, this state of affairs closely resembles the state of quantum field theory (QFT) prior to Wilson’s seminal work on the renormalization group [35, 36]. At that time, QFT was a wildly powerful, but largely unwieldy tool. Wilsonian RG gave a methodology for organizing our thinking about QFTs which in turn has given us more control over how we can use and study them. Our hope is that the approach we have outlined in this paper can serve a similar organizing role for machine learning and the study of NNs.
@@ Acknowledgements
We thank Pietro Rotondo and Veronica Guidetti for their helpful comments on the draft. We also wish to thank Jim Halverson for useful discussions on the NNFT correspondence at StringData 2023 and David Berman for his insight and many useful discussions over the course of this work. Finally, we thank Edward Hirst, Nathaniel Craig, Katy Craig, Emanuele Gendy, Ro Jefferson, and Zohar Ringel for useful discussions.